What is the probability that you guess the correct 4 digit pin of a bank account if you know it repeats two numbers at least once?
Answer #1

here total number of possible outcomes =10^4 =10000 (since 10 choices for each number)

number of ways all 4 digits are distinct =10*9*8*7 =5040

therefore number of ways t repeats two numbers at least once =10000-5040=4960

hence probability =4960/10000 =0.4960